FTMW/CVI: FOURIER TRANSFORM MICROWAVE / COHERENCE SPECTROSCOPY WITH VIRTUAL INSTRUMENTS

\underline{JENS-UWE GRABOW}, Lehrgebiet Physikalische Chemie A, Institut für Physikalische Chemie und Elektrochemie, Callinstrasse 3-3a, Universtät Hannover, D-30167 Hannover.

Modern Fourier transform microwave (FTMW) spectroscopy, first implemented as an impulse excitation technique in a steady gas-waveguide apparatus^{a}, started only a quarter of a century ago. Soon after, the method was supplemented for special applications, e.g. Stark-effect^{b}, double resonance^{c}, and 2D^{d} spectroscopy. The subsequent implementation of the technique in a supersonic jet-resonator apparatus^{e} was followed by similar developments, i.e. Stark-effect^{f}, double resonance^{g}, laser ablation^{h}, and DC-discharge^{i}. Automation^{j} along with the “coaxially oriented beam-resonator arrangement”^{k} (COBRA) has greatly increased efficiency, resolution, and sensitivity of current FTMW spectrometers. We present the implementation of a complete COBRA-FTMW spectrometer featuring the capabilities given above. The entire experiment is operated, either interactivle or automatically, by a graphical user interface (GUI)-based program which is capable of waveguide and resonator applications. Developed in the C programming language, while following the concept of virtual instruments (VI), the software can co-operate with a wide variety of PCI and IEEE hardware components to build the instrument. Only one application specific logic circuit, which will also be presented, is needed.

aJ. Ekkers and W. H. Flygare, Rev. Sci. Instrum. 47, 448(1976).

bG. Bestmann and H. Dreizler, Z. Naturforsch. 37a, 615(1982).

cH. Dreizler, E. Fliege, H. Mäder, and W. Stahl, Z. Naturforsch. 37a, 1266(1982).

dD. A. Andrews and J. G. Baker, J. Phys. B20, 5705(1987).

eT. J. Balle and W. H. Flygare, Rev. Sci. Instrum. 52, 33(1981).

fL. H. Coudert, F. J. Lovas, R. D. Suenram, and J. T. Hougen, J. Chem. Phys. 87, 6290(1987).

gL. Martinache, S. Jans-Bürli, B. Vogelsanger, W. Kresa, and A. Bauder, Chem. Phys. Lett. 149, 424(1988).

hR. D. Suenram, F. J. Lovas, and K. Matsumura, Ap. J. Lett. 342, L103(1989).

iJ.-U. Grabow, N. Heineking, and W. Stahl, Z. Naturforsch. 46a, 914(1991).

jU. Andresen, H. Dreizler, J.-U. Grabow and W. Stahl, Rev. Sci. Instrum. 61, 3694(1990).

kJ.-U. Grabow and W. Stahl, Z. Naturforsch. 45a, 1043(1990)